OpenStudy (anonymous):

If I have 3 red marbles and 7 blue marbles and pick and replace marbles with like marbles, what is the probability of picking 3 red marbles in a row?

OpenStudy (anonymous):

\[\Big\lgroup\frac{Num_Red}{Total_Marbles}\Big\rgroup*\Big\lgroup\frac{Num_Red}{Total_Marbles}\Big\rgroup*\Big\lgroup\frac{Num_Red}{Total_Marbles}\Big\rgroup\]

OpenStudy (anonymous):

Imagine this as 3 independent trials using 3 bags of marbles in a row. Each bag has 3 of 10 marbles that are red. The result of one bag is independent of the result of the other bags -- so there is the same fraction, multiplied 3 times for the three seperate bags.